# Recurrent Inference Machine for Medical Image Registration

### v1

@@ Abstract
Image registration is essential for medical image applications where alignment of voxels across multiple images is needed for qualitative or quantitative analysis. With recent advancements in deep neural networks and parallel computing, deep learning-based medical image registration methods become competitive with their flexible modelling and fast inference capabilities. However, compared to traditional optimization-based registration methods, the speed advantage may come at the cost of registration performance at inference time. Besides, deep neural networks ideally demand large training datasets while optimization-based methods are training-free. To improve registration accuracy and data efficiency, we propose a novel image registration method, termed Recurrent Inference Image Registration (RIIR) network. RIIR is formulated as a meta-learning solver to the registration problem in an iterative manner. RIIR addresses the accuracy and data efficiency issues, by learning the update rule of optimization, with implicit regularization combined with explicit gradient input.
We evaluated RIIR extensively on brain MRI and quantitative cardiac MRI datasets, in terms of both registration accuracy and training data efficiency. Our experiments showed that RIIR outperformed a range of deep learning-based methods, even with only $5\%$ of the training data, demonstrating high data efficiency. Key findings from our ablation studies highlighted the important added value of the hidden states introduced in the recurrent inference framework for meta-learning. Our proposed RIIR offers a highly data-efficient framework for deep learning-based medical image registration.
Keywords: Deep Learning, Image Registration, Recurrent Inference Machine
1 Introduction
@@ 1.1.2 Iterative Deep Learning-based Registration
However, a one-step inference strategy may struggle when predicting large and complex transformations (Hering et al., 2019; Zhao et al., 2019). In contrast to one-step deep learning-based registration methods, recent work adopted iterative processes, reincarnating the conventional pipeline of optimization for medical image registration, either in terms of image resolution (Hering et al., 2019; Fechter and Baltas, 2020; Xu et al., 2021; Liu et al., 2021), multiple optimization steps (Zhao et al., 2019; Sandkühler et al., 2019; Kanter and Lellmann, 2022), or combined (Qiu et al., 2022). In Sandkühler et al. (2019), the use of RNN with gated recurrent unit (GRU) (Chung et al., 2014) was considered, where each step progressively updates the transformation by adding an independent parameterized transformation. Another multi-step method proposed in Zhao et al. (2019) uses recursive cascaded networks to generate a sequence of transformations, which is then composed to get the final transformation. However, the method requires independent modules for each step, which can be memory-inefficient.Hering et al. (2019) proposed a variational method on different levels of resolution, where the final transformation is the composition of the transformations from coarse- to fine-grained. Fechter and Baltas (2020) addresses the importance of data efficiency of deep learning-based models by evaluating the model performance when data availability is limited, and a large domain shift exists. A more recent work proposed in Qiu et al. (2022), Gradient Descent Network for Image Registration (GraDIRN), integrates multi-step and multi-resolution for medical image registration. Specifically, the update rule follows the idea of conventional optimization by deriving the gradient of the similarity term w.r.t. the current transformation and using a CNN to estimate the gradient of the regularization term. Though the direct influence of the gradient term shows to be minor compared to the CNN output (Qiu et al., 2022), the method bridges gradient-based optimization and deep learning-based methods. The method proposed in Kanter and Lellmann (2022) used individual long short-term memory (LSTM) modules for implementing recurrent refinement of the transformation. However, the scope of the work is limited to affine transformation, which only serves as an initialization for the conventional medical image registration pipeline.
@@ 1.1.3 Meta-Learning and Recurrent Inference Machine
In the spirit of meta-learning, RIM was developed by Putzky and Welling (2017) to solve inverse problems. RIM learns a single recurrent architecture that shares the parameters across all iterations, with internal states passing through iterations (Putzky and Welling, 2017). In the context of meta-learning, RIM distinguishes two tasks of different levels: the ‘inner task’, which focuses on solving a specific inverse problem (e.g., superresolution of an image), and the ‘outer task’, aimed at optimizing the optimization process itself. This setting enables RIM to efficiently learn and apply optimization strategies to complex problems. RIM has shown robust and competitive performance across different application domains, from cosmology (Morningstar et al., 2019; Modi et al., 2021) to medical imaging (Karkalousos et al., 2022; Lønning et al., 2019; Putzky et al., 2019; Sabidussi et al., 2021, 2023). They all aim to solve an inverse problem with a known differentiable forward model in closed form, such as Fourier transform with sensitivity map and sampling mask in MRI reconstruction (Lønning et al., 2019).
However, the definition of an explicit forward model does not exist for the medical image registration task. In this work, we sought to extend the framework of RIM, which demonstrated state-of-the-art performance in medical image reconstruction challenges (Muckley et al., 2021; Putzky et al., 2019; Zbontar et al., 2018), to the medical image registration problem. The same formulation can be generalized to other high-dimensional optimization problems where explicit forward models are absent but differentiable evaluation metrics are available.
@@ 2.1 Deformable Image Registration
Deformable image registration aims to align a moving image $I_{\text{mov}}$ to a fixed image $I_{\text{fix}}$ by determining a transformation $\bm{\phi}$ acting on the shared coordinates $\bm{x}$, such that the transformed image $I_{\text{mov}}\circ\bm{\bm{\phi}}$ is similar enough to $I_{\text{fix}}$. The similarity is often evaluated by a scalar-valued metric. In deformable image registration, $\bm{\bm{\phi}}$ is considered to be a relatively small displacement added to the original coordinate $\bm{x}$, expressed as $\bm{\phi}=\bm{x}+u(\bm{x})$. Since the transformation $\bm{\phi}$ is calculated between the pair $(I_{\text{mov}},I_{\text{fix}})$, the process is often referred to as pairwise registration (Balakrishnan et al., 2019). Finding such transformation $\bm{\phi}$ in pairwise registration can be viewed as the following optimization problem:
@@ 2.2 Recurrent Inference Machine (RIM)
where $g_{\theta}$ is a neural network parameterized by $\theta$. In RIM, the prior distribution $p_{\text{prior}}(\bm{x})$ is implicitly integrated into the parameterized neural network $g_{\theta}$ which is trained with a weighted sum of the individual prediction losses between $\bm{x}$ and $\bm{x}_{t}$ (e.g., the mean squared loss) at each time step $t$.
@@ 2.3 Recurrent Inference Image Registration Network (RIIR)
Inspired by the formulation of RIM and the optimization nature of medical image registration, we present a novel deep learning-based image registration framework, named the Recurrent Inference Image Registration Network (RIIR). The overview of our proposed framework can be found in Fig. 1. The proposed framework performs an end-to-end iterative prediction of a dense transformation $\bm{\phi}$ in $T$ steps for pairwise registration: Given the input image pair $\left(I_{\text{mov}},I_{\text{fix}}\right)$, the optimization problem in Eq. 1 can be solved by the iterative update of $\bm{\phi}$. And the update rule at step $t\in\{0,1,\ldots,T-1\}$ is:
@@
where $\bm{\phi}_{0}$ is initialized as an identity mapping $\bm{\phi}_{0}(\bm{x})=\bm{x}$. The update at step $t$, $\Delta\bm{\phi}_{t}$, is calculated by a recurrent update network $g_{\theta}$ by taking a channel-wise concatenation of
$\{\bm{\phi}_{t},\nabla_{\bm{\phi}_{t}}\mathcal{L}_{\text{inner}}\left(I_{\text{mov}}\circ\bm{\phi}_{t},I_{\text{fix}}\right)\}$ (9)
@@
Originally, RIM aimed to learn a recurrent solver for an inverse problem where the forward model from signal to measurement is known, such as quantitative mapping (Sabidussi et al., 2021) or MRI reconstruction (Lønning et al., 2019). However, the optimization of image registration is in an unsupervised manner with the forward model from the measurement pair $(I_{\text{mov}},I_{\text{fix}})$ to the deformation field $\bm{\phi}$ unknown. In this work, we extend the framework of RIM to be amendable to a broader range of tasks including image registration. We design the inner loss $\mathcal{L}_{\text{inner}}$ by adapting the optimization objective in Eq. 1. Specifically, we use the similarity part $\mathcal{L}_{\text{sim}}$ in Eq. 1 as the inner loss at time step $t$:
The explicit modelling of $\mathcal{L}_{\text{inner}}$ also helps the auto differentiation for calculating its gradient while making the regularization part to be learned implicitly in $g_{\theta}$ during the training.
In the implementation of RIM, the iterative update Eq. 8 is achieved by a recurrent neural network (RNN) to generalize the update rule in Eq. 5 with hidden memory state variable $\bm{h}$ estimated for each time step $t$. Unlike previous RIM-based works (Putzky et al., 2019; Sabidussi et al., 2021) which use two linear gated recurrent units (GRU) to calculate the hidden states $\bm{h}_{t}$, in RIIR, two convolutional gated recurrent units (ConvGRU) (Shi et al., 2015) are used to better preserve spatial correlation in the image. We further investigate the necessity of including such two-level recurrent structures in our experiment, particularly considering potential complexities in constructing computation graphs for neural networks. The iterative update equations of RIIR at step $t$ have the following form, with the hidden memory states:
$\left\{\Delta\bm{\phi}_{t},\bm{h}_{t+1}\right\}$ $=g_{\theta}(\bm{\phi}_{t},\nabla_{\bm{\phi}_{t}}\mathcal{L}_{\text{inner}}\left(I_{\text{mov}}\circ\bm{\phi}_{t},I_{\text{fix}}\right),\bm{h}_{t}),$ (11)
@@
where $\bm{h}_{t}=\{\bm{h}_{t}^{1},\bm{h}_{t}^{2}\}$ denotes the two-level hidden memory states at step $t$. The size of $\bm{h}_{t}$ depends on the size of input image pair $(I_{\text{mov}},I_{\text{fix}})$ with multiple channels. For $t=1$, $\bm{h}_{1}$ is initialized to a zero input. We name our network $g_{\theta}$ as RIIR Cell, with its detailed architecture illustrated in Fig. 1. To address the difference between our RIIR from the existing gradient-based iterative algorithm (GraDIRN) (Qiu et al., 2022) under the same definition of $\mathcal{L}_{\text{inner}}$ as in Eq. 10, RIIR uses the gradient of inner loss as the neural network input to calculate the incremental update. On the other hand, GraDIRN takes the channel-wise warped image pair $(I_{\text{mov}}\circ\phi,I_{\text{fix}})$ and deformation field $\phi$ as the input to the network to output regularization update in Eq. 1, while the gradient of $\mathcal{L}_{\text{inner}}$ is added to the update without any further processing. Since the network $g_{\theta}$ learns an incremental update with the gradient of inner loss as compressed information, the training of RIIR can be more efficient, which will be discussed in the experiment section.
Unlike previous work in deep learning-based iterative deformable image registration methods which does not incorporate internal hidden states (Zhao et al., 2019; Fechter and Baltas, 2020; Qiu et al., 2022), we propose to combine the gradient information and hidden states as the network input. Using $\bm{h}_{t}$ also suggests an analogy with gradient-based optimization methods such as the Limited-memory Broyden–Fletcher–Goldfarb–Shanno algorithm (L-BFGS) to track and memorize progression (Putzky and Welling, 2017). To substantiate this design, the input selections of RIIR will be further ablation-studied and discussed in our experiments.
@@
where $w_{t}$ is a (positive) scalar indicating the weight of step $t$. In our experiment, both uniform ($w_{t}=\frac{1}{T}$) and exponential weights ($w_{t}=10^{\frac{t-1}{T-1}}$) are considered and will be compared in the experiments. It is noticeable that the design of using (weighted) average of step-wise loss also makes our proposed RIIR different from other iterative deep learning-based methods (Qiu et al., 2022; Zhao et al., 2019), addressing the fact that early steps in the prediction process were neglected before.

### v2

@@ Abstract
Image registration is essential for medical image applications where alignment of voxels across multiple images is needed for qualitative or quantitative analysis. With recent advances in deep neural networks and parallel computing, deep learning-based medical image registration methods become competitive with their flexible modelling and fast inference capabilities. However, compared to traditional optimization-based registration methods, the speed advantage may come at the cost of registration performance at inference time. Besides, deep neural networks ideally demand large training datasets while optimization-based methods are training-free. To improve registration accuracy and data efficiency, we propose a novel image registration method, termed Recurrent Inference Image Registration (RIIR) network. RIIR is formulated as a meta-learning solver for the registration problem in an iterative manner. RIIR addresses the accuracy and data efficiency issues, by learning the update rule of optimization, with implicit regularization combined with explicit gradient input.
We extensively evaluated RIIR on brain MRI, lung CT, and quantitative cardiac MRI datasets, in terms of both registration accuracy and training data efficiency. Our experiments showed that RIIR outperformed a range of deep learning-based methods, even with only $5\%$ of the training data, demonstrating high data efficiency. Key findings from our ablation studies highlighted the important added value of the hidden states introduced in the recurrent inference framework for meta-learning. Our proposed RIIR offers a highly data-efficient framework for deep learning-based medical image registration.
Keywords: Image Registration, Deep Learning, Recurrent Inference Machine
1 Introduction and Related Works
@@ 1.1.2 Iterative Deep Learning-based Registration
However, a one-step inference strategy may struggle when predicting large and complex transformations (Hering et al., 2019; Zhao et al., 2019). In contrast to one-step deep learning-based registration methods, recent work adopted iterative processes, reincarnating the conventional pipeline of optimization for medical image registration, either in terms of image resolution (Hering et al., 2019; Mok and Chung, 2020; Fechter and Baltas, 2020; Xu et al., 2021; Liu et al., 2021), multiple optimization steps (Zhao et al., 2019; Sandkühler et al., 2019; Falta et al., 2022; Kanter and Lellmann, 2022), or combined (Qiu et al., 2022). In Sandkühler et al. (2019), the use of RNN with gated recurrent unit (GRU) (Chung et al., 2014) was considered, where each step progressively updates the transformation by adding an independent parameterized transformation. Another multi-step method proposed in Zhao et al. (2019) uses recursive cascaded networks to generate a sequence of transformations, which is then composed to get the final transformation. However, the method requires independent modules for each step, which can be memory-inefficient.Hering et al. (2019) proposed a variational method on different levels of resolution, where the final transformation is the composition of the transformations from coarse- to fine-grained. Fechter and Baltas (2020) addresses the importance of data efficiency of deep learning-based models by evaluating the model performance when data availability is limited, and a large domain shift exists. Falta et al. (2022) proposed an iterative method named Learn-to-Optimise (L2O) to emulate the gradient-based optimization in lung CT registration. Unlike the fully unsupervised training scheme, the method utilizes a deep supervision strategy on the generated key points with a recurrent use of U-Net. Noticeably, the method uses additional input feature modalities including dynamically sampled coordinates and MIND features (Heinrich et al., 2012) to enhance the model. A more recent work proposed in Qiu et al. (2022), Gradient Descent Network for Image Registration (GraDIRN), integrates multi-step and multi-resolution for medical image registration. Specifically, the update rule follows the idea of conventional optimization by deriving the gradient of the similarity term w.r.t. the current transformation and using a CNN to estimate the gradient of the regularization term. Though the direct influence of the gradient term shows to be minor compared to the CNN output (Qiu et al., 2022), the method bridges gradient-based optimization and deep learning-based methods. The method proposed in Kanter and Lellmann (2022) used individual long short-term memory (LSTM) modules for implementing recurrent refinement of the transformation. However, the scope of the work is limited to affine transformation, which only serves as an initialization for the conventional medical image registration pipeline.
@@ 1.1.3 Meta-Learning and Recurrent Inference Machine
In the spirit of meta-learning, RIM was developed by Putzky and Welling (2017) to solve inverse problems. RIM learns a single recurrent architecture that shares the parameters across all iterations, with internal states passing through iterations (Putzky and Welling, 2017). In the context of meta-learning, RIM distinguishes two tasks of different levels: the ‘inner task’, which focuses on solving a specific inverse problem (e.g., superresolution of an image), and the ‘outer task’, aimed at optimizing the optimization process itself. This setting enables RIM to efficiently learn and apply optimization strategies to complex problems. Therefore, RIM only has one neural network component which learns the outer task. RIM has shown robust and competitive performance across different application domains, from cosmology (Morningstar et al., 2019; Modi et al., 2021) to medical imaging (Karkalousos et al., 2022; Lønning et al., 2019; Putzky et al., 2019; Sabidussi et al., 2021, 2023). To the best of our knowledge, most applications of RIM aim to solve an inverse problem with a known differentiable forward model in closed form, such as Fourier transform with sensitivity map and sampling mask in MRI reconstruction (Lønning et al., 2019).
However, the definition of an explicit forward model does not exist for the medical image registration task. Although RIM does not require a forward model by its design, the absence of a concrete forward model makes the problem more complicated. In this case, our formulation is similar to a realization of iterative amortized inference (Marino et al., 2018). In Marino et al. (2018), a variational auto-encoder (VAE) framework is studied to learn the amortized optimization process given the input data and approximate posterior gradients where the likelihood under certain forward model could be absent. In this work, we sought to extend the framework of RIM, which demonstrated state-of-the-art performance in medical image reconstruction challenges (Muckley et al., 2021; Putzky et al., 2019; Zbontar et al., 2018), to the medical image registration problem which relaxes the need of gradient likelihood under specific forward model solely. The same formulation can be generalized to other high-dimensional optimization problems where explicit forward models are absent but differentiable evaluation metrics are available.
@@ 2.1 Deformable Image Registration
Deformable image registration aims to align a moving image $I_{\text{mov}}$ to a fixed image $I_{\text{fix}}$ by determining a transformation $\bm{\phi}$ acting on the shared coordinates $\bm{\chi}$, such that the transformed image $I_{\text{mov}}\circ\bm{\bm{\phi}}$ is similar enough to $I_{\text{fix}}$. The similarity is often evaluated by a scalar-valued metric. In deformable image registration, $\bm{\bm{\phi}}$ is considered to be a relatively small displacement added to the original coordinate $\bm{\chi}$, expressed as $\bm{\phi}=\bm{\chi}+u(\bm{\chi})$. Since the transformation $\bm{\phi}$ is calculated between the pair $(I_{\text{mov}},I_{\text{fix}})$, the process is often referred to as pairwise registration (Balakrishnan et al., 2019). Finding such transformation $\bm{\phi}$ in pairwise registration can be viewed as the following optimization problem:
@@ 2.2 Recurrent Inference Machine (RIM)
where $g_{\theta}$ is a neural network parameterized by $\theta$. In RIM, the prior distribution on the data prior (regularization) $p_{\text{prior}}(\bm{x})$ is implicitly integrated into the parameterized neural network $g_{\theta}$ which is trained with a weighted sum of the individual prediction losses between $\bm{x}$ and $\bm{x}_{t}$ (e.g., the mean squared loss) at each time step $t$.
@@ 2.3 Recurrent Inference Image Registration Network (RIIR)
Inspired by the formulation of RIM and the optimization nature of medical image registration, we present a novel deep learning-based image registration framework, named the Recurrent Inference Image Registration Network (RIIR). The overview of our proposed framework can be found in Fig. 1.
Originally, RIM aimed to learn a recurrent solver for an inverse problem where the forward model from signal to measurement is known for inverse problems, such as quantitative mapping (Sabidussi et al., 2021) or MRI reconstruction (Lønning et al., 2019). Similarly to RIM in other medical image applications, the regularization is proposed to be learned implicitly in the neural network. Therefore, we determine the inner loss $\mathcal{L}_{\text{inner}}$ by adapting the optimization objective in Eq. 1. Specifically, we use the similarity part $\mathcal{L}_{\text{sim}}$ in Eq. 1 as the inner loss at time step $t$:
The gradient of $\mathcal{L}_{\text{inner}}$ can be calculated using auto differentiation.
Nevertheless, in the standard RIM framework on the optimization objective in Eq. 1, the displacement fields serve as the primary optimization signal which may overlook the rich structural and contextual information present in the original and warped images. This information loss can be critical, particularly for capturing implicit regularization introduced by the structures in original images, which directly influences $\mathcal{L}_{\text{outer}}$ and the goal of optimizing the evaluation metric (e.g., the overlapping of tissues or organs) which is usually not equivalent to $\mathcal{L}_{\text{outer}}$. We recognize such potential loss of information when relying solely on the gradient of the inner loss w.r.t. to $\bm{\phi}$. To address this, we extend the RIM framework, drawing inspiration from a more generalized formulation: the iterative amortized inference (Marino et al., 2018), ensuring that the model leverages the original images more effectively in the iterative optimization process. Such an extension takes the warped images as input for optimization, while not changing the optimization objective but enriching the information that could be passed into the neural network.
With this modification, our proposed framework performs an end-to-end iterative prediction of a dense transformation $\bm{\phi}$ in $T$ steps for pairwise registration: Given the input image pair $\left(I_{\text{mov}},I_{\text{fix}}\right)$, the optimization problem in Eq. 1 can be solved by the iterative update of $\bm{\phi}$. And the update rule at step $t\in\{0,1,\ldots,T-1\}$ is:
@@
where $\bm{\phi}_{0}$ is initialized as an identity mapping $\bm{\phi}_{0}(\bm{\chi})=\bm{\chi}$. The update at step $t$, $\Delta\bm{\phi}_{t}$, is calculated by a recurrent update network $g_{\theta}$ by taking a channel-wise concatenation of
$\{\bm{\phi}_{t},\nabla_{\bm{\phi}_{t}}\mathcal{L}_{\text{inner}}\left(I_{\text{mov}}\circ\bm{\phi}_{t},I_{\text{fix}}\right),I_{\text{mov}}\circ\bm{\phi}_{t},I_{\text{fix}}\}$ (10)
@@
In the implementation of RIM, the iterative update Eq. 9 is achieved by a recurrent neural network (RNN) to generalize the update rule in Eq. 5 with hidden memory state variable $\bm{h}$ estimated for each time step $t$, which is the only neural network component in the whole pipeline. Unlike previous RIM-based works (Putzky et al., 2019; Sabidussi et al., 2021) which use two linear gated recurrent units (GRU) to calculate the hidden states $\bm{h}_{t}$, in RIIR, two convolutional gated recurrent units (ConvGRU) (Shi et al., 2015) are used to better preserve spatial correlation in the image. We further investigate the necessity of including such two-level recurrent structures in our experiment, particularly considering potential complexities in constructing computation graphs for neural networks. The iterative update equations of RIIR at step $t$ have the following form, with the hidden memory states:
$\left\{\Delta\bm{\phi}_{t},\bm{h}_{t+1}\right\}$ $=g_{\theta}(\bm{\phi}_{t},\nabla_{\bm{\phi}_{t}}\mathcal{L}_{\text{inner}}\left(I_{\text{mov}}\circ\bm{\phi}_{t},I_{\text{fix}}\right),I_{\text{mov}}\circ\bm{\phi}_{t},I_{\text{fix}},\bm{h}_{t}),$ (11)
@@
where $\bm{h}_{t}=\{\bm{h}_{t}^{1},\bm{h}_{t}^{2}\}$ denotes the two-level hidden memory states at step $t$. The size of $\bm{h}_{t}$ depends on the size of input image pair $(I_{\text{mov}},I_{\text{fix}})$ with multiple channels. For $t=1$, $\bm{h}_{1}$ is initialized to a zero input. We name our network $g_{\theta}$ as RIIR Cell, with its detailed architecture illustrated in Fig. 1. To address the difference between our RIIR from the existing gradient-based iterative algorithm (GraDIRN) (Qiu et al., 2022) under the same definition of $\mathcal{L}_{\text{inner}}$ as in Eq. 8, RIIR uses the gradient of inner loss as the neural network input to calculate the incremental update. On the other hand, GraDIRN takes the channel-wise warped image pair $(I_{\text{mov}}\circ\phi,I_{\text{fix}})$ and deformation field $\phi$ as the input to the network to output regularization update in Eq. 1, while the gradient of $\mathcal{L}_{\text{inner}}$ is added to the update of the deformation without any further processing thus does not directly affect the network part of the pipeline.
Unlike previous work in deep learning-based iterative deformable image registration methods which does not incorporate internal hidden states (Zhao et al., 2019; Fechter and Baltas, 2020; Qiu et al., 2022), we propose to combine the gradient information and hidden states as the network input. Our method also differs from Falta et al. (2022) in several aspects: in Falta et al. (2022), the input consists of a collection of images, displacement, sampled coordinates and MIND features; also, the part of the U-Net output channels serve as hidden states, instead of using dedicated ConvGRU units as in our design. Using $\bm{h}_{t}$ also suggests an analogy with gradient-based optimization methods such as the Limited-memory Broyden–Fletcher–Goldfarb–Shanno algorithm (L-BFGS) to track and memorize progression (Putzky and Welling, 2017). To substantiate this design, the input selections of RIIR will be further ablation-studied and discussed in our experiments.
@@
where $w_{t}$ is a (positive) scalar indicating the weight of step $t$. In our experiment, both uniform ($w_{t}=\frac{1}{T}$) and exponential weights ($w_{t}=10^{\frac{t-1}{T-1}}$) are considered and will be compared in the experiments. It is noticeable that the design of using a (weighted) average of the stepwise loss also makes our proposed RIIR different from other iterative deep learning-based methods (Qiu et al., 2022; Zhao et al., 2019) which use only the final output to calculate the loss, and Falta et al. (2022) use a uniform weight, addressing the fact that early steps in the prediction process were neglected before.